# Customizing Spider Silk: Generative Models with Mechanical Property Conditioning for Protein Engineering

### v1

@@ 4.4 MaSp Motif-Property Correlation Analysis
The relationship between MaSp motifs and mechanical properties in spider dragline silk offers critical insights into how specific protein sequence patterns influence silk performance. Recent high-throughput studies have quantified key mechanical metrics -including toughness, Young’s modulus, tensile strength, and strain at break - of dragline silk from various spider species, identifying correlations between motifs in the repetitive core region of MaSps and variations in these properties [15].
@@
In previous studies, correlation analyses have primarily focused on the distinct motifs present in different MaSp types (MaSp1, MaSp2, and MaSp3) and their specific influences on mechanical properties [25, 49, 50, 15]. By contrast, the current work considers MaSp more abstractly, providing an overarching view of how various motifs may affect mechanical performance.
@@
Figure 4 presents the correlation heatmap between sequential features and mechanical properties. The sequential features include the count and coverage (motif size×countsequence $\frac{\text{motif size}\times\text{count}}{\text{sequence length}}$) of the specified motifs. The heatmap indicates that correlation values remain relatively low. While no strong correlations were observed, this visualization is still informative, demonstrating that individual motifs alone cannot fully explain the relationship between sequence variation and mechanical properties. This highlights the intricate and multi-factorial nature of silk mechanics, where the combined influence of multiple motifs and structural contexts collectively shapes the observed mechanical behavior.
@@
In the following we mention the row vise analysis of the heatmap:
@@
Toughness is the total energy absorbed before failure. Higher toughness in dragline silk is associated with motifs that enhance tensile strength and/or extensibility. The correlation heatmap reveals that YGQGG count (+0.198) and coverage (+0.188) both show moderate positive correlations with toughness. Poly-Ala coverage also correlates positively with toughness, though slightly less strongly (+0.135). The SV count showed negative correlation with toughness (-0.113).
@@
Young’s modulus represents the initial stiffness of silk fibers and is influenced by motifs that foster $\beta$-sheet crystallinity. For example, poly-Ala count correlates positively with Young’s modulus (+0.067). In our analysis, we also observed a positive impact of the count of YGQGG motif on Young’s modulus (+0.085), whereas the count of AGQG (-0.033) and SV (-0.126) motifs showed negative correlations.
@@
Tensile strength is the maximum stress that the silk fiber can withstand before failure. The correlation heatmap indicates that many of the same motifs influencing toughness also affect strength. The YGQGG motif count is positively correlated with tensile strength (+0.206) [15], making it one of the strongest sequence predictors of a stronger fiber in our analysis. Poly-Ala motifs also positively correlate with tensile strength (+0.107).
@@
Strain at break is the maximal elongation (expressed as a fraction of its original length) that the silk fiber can achieve before breaking. The correlations highlight AGQG coverage (–0.103) is the most striking correlation overall (and the only moderately strong negative value). Other features in this row hover near zero, indicating minimal correlation. The results also show impact of QQ motif on strain at break. Other features in this row hover near zero, indicating minimal or no correlation.
In summary, the correlation between mechanical properties and individual motifs in the generated sequences follows patterns reported in previous studies [15, 26]. However, the overall weak correlations suggest that individual motifs alone cannot account for the observed variations in fiber mechanical properties [15]. This study builds upon existing sequence-property relationships and leverages machine learning to better capture the complex interplay between sequential motifs and mechanical properties, enabling the design of novel MaSp sequences.
@@ 5.1.1 Conditional Generation Assessment: Evaluation on test dataset
To assess the self-consistency of the trained model, we curated a test set of 20 instances, sampled from the original 592-instance Spider Silkome dataset, which consists of MaSp repeats along with their corresponding mechanical properties. The details of the test set are mentioned in the supporting document SD2. The mechanical properties from the test set were used for conditional sequence generation, allowing for a direct comparison between the generated and natural sequences.
To evaluate the model’s ability to generate sequences that correspond to a known set of mechanical properties, we analyzed the test set by comparing the generated MaSp repeats with their original counterparts. The evaluation results, presented in Figure 5, assess the generated sequences at both the sequence level and the structural level, providing insights into how well the model captures the underlying patterns of spider silk proteins.
@@ 5.1.2 Novelty Assessment: BLAST Evaluation
To evaluate the novelty of the generated MaSp repeats, we performed BLAST analyses against a broader Spider Silkome spidroin repeat database. The search space consists of 11K naturally occurring spidroin sequences. It includes seven primary types of spider silk proteins major ampullate spidroins (MaSp), flagelliform spidroins (Flag), minor ampullate spidroins (MiSp), aggregate spidroins (AgSp), pyriform spidroins (PySp), aciniform spidroins (AcSp), and cylindriform/tubuliform spidroins (CySp) [51]. A subset of seven generated sequences was compared with the two most closely related natural sequences, designated BLAST1 (top match) and BLAST2 (second top match) based on BLAST results. The selection was based on high query coverage, percent identity, and sequence length similarity to the generated sequences. This dataset is referred to as the "BLAST set", with details of all sequences and their sources provided in the supporting document SD3. All selected BLAST matches had an expect value (E) of $E\leq e^{-10}$, which ensured statistically significant similarities. Furthermore, we selected matches spanning different MaSp subtypes (MaSp1, MaSp2, MaSp3) to provide a comprehensive comparative analysis. The expect value (E) quantifies the probability of obtaining a match by chance in a given database size, decreasing exponentially as the alignment score (S) increases.
To assess uniqueness, we follow established criteria where sequences with similarity values below 50–60% are considered novel [52]. The generated sequences meet this threshold when compared to broader protein databases, demonstrating the model’s ability to design distinct sequences that either do not exist in nature or have not yet been observed. Additionally, BLAST search results consistently classify the most similar existing sequences as MaSp, aligning with the intended spidroin type. This indicates that the generative modeling approach effectively captures the defining characteristics of specific spider silk proteins.
@@ 5.2 Evaluation of Mechanical Property Prediction
Table 1: Comparison of Generated and Reference Properties
Metric Value
Pearson Correlation ($r$) 0.3911
Mean Absolute Error (MAE) 0.0587
Root Mean Square Error (RMSE) 0.0716
Cosine Similarity 0.9465
Figure 7: Comparison of predicted and original mechanical properties for MaSp repeats in the test set. Scatter plots illustrate the relationship between original and predicted values for toughness, Young’s modulus, tensile strength, and strain at break. Property values are normalized between 0 and 1 with each plot displaying the corresponding Pearson correlation coefficient (r), which ranges from -1 to +1. The red dashed diagonal line represents the ideal 1:1 correlation, indicating perfect agreement between predicted and actual values. The observed trends suggest a significant correlation between generated and original mechanical properties, demonstrating the effectiveness of the prediction model.
To evaluate models’ capability to estimate mechanical properties for a given MaSp repeat, we use the test set of 20 instances which we mentioned above. For each sequence in the test set, we run inference for the reverse task of the model and generate mechanical properties. The similarity between the generated and reference properties was assessed using four key statistical metrics: Pearson’s correlation coefficient (r), mean absolute error (MAE), root mean square error (RMSE), and cosine similarity, as detailed in Table 1. These metrics provide quantitative insight into the degree to which the generated sequences replicate the structural and mechanical trends observed in the reference dataset.
The Pearson Correlation Coefficient ($r$) was found to be 0.3911, indicating a moderate positive correlation between the generated and reference properties. Although this suggests that the generated sequences capture some variations present in the reference data, the relationship is not strictly linear. The Mean Absolute Error (MAE) of 0.0587 represents the average absolute difference between generated and reference values. The relatively low MAE suggests that the generated properties closely approximate the reference values, with minor deviations. The Root Mean Square Error (RMSE) of 0.0716 is slightly higher than the MAE, reflecting its sensitivity to larger deviations. This indicates that while most generated values align well with reference values, occasional discrepancies exist but remain within an acceptable range. Finally, the Cosine Similarity of 0.9465 demonstrates a high degree of directional similarity between the generated and reference properties. This suggests that while absolute values may not always align perfectly, the generated sequences effectively maintain overall structural and mechanical trends.
A property-level analysis is presented in Figure 7, further evaluating the accuracy of the estimated mechanical properties. Figure 7 (a) illustrates the correlation between original and generated values across all four properties, reaffirming the model’s capability to estimate mechanical properties with reasonable accuracy. Overall, these findings highlight the generative model’s effectiveness in capturing key mechanical characteristics, making it a reliable tool for designing bio-inspired materials.
@@ 5.3.2 Without first level fine tuning
ProtGPT2 is a decoder-only transformer model that has been pre-trained on the protein space using the UniRef50 database (version 2021_04) [21], which contains 10 million protein instances. However, the Sequence-to-Mechanical Property Correlation dataset, curated from the Spider Silkome database, contains only 592 instances. This small dataset is insufficient for training the model to generate MaSp repeat sequences while simultaneously learning the correlation between sequences and mechanical properties. To address this limitation, we split the training into two stages.
@@
To further investigate these discrepancies, we compare the structural predictions of sequences generated by a model trained with only one fine-tuning stage against those produced by a model trained with both levels of fine-tuning. Figure 9 illustrates the molecular configurations of the generated sequences, revealing structural differences arising from the absence of MaSp-specific fine-tuning. Structural predictions were obtained using Omegafold [53], providing insights into the conformational tendencies of the sequences. Notably, the generated sequences without level 1 fine-tuning exhibited helical structures, which are typically associated with non repeat regions like terminal domains (NTD/CTD). These unwanted elements further highlight the necessity of level 1 fine-tuning in ensuring the correct generation of MaSp repeats.
@@
Here, we analyze the effect of skipping level 1 fine-tuning on the reverse task of mechanical property estimation. To evaluate performance, we used the same test set of 20 instances previously employed to assess the proposed pipeline. Specifically, we provided the sequence generation prompts from the test set and compared the predicted properties from both model variants—one with level 1 fine-tuning and one without.
Table 3: Comparison of Generated and Reference Properties
Metric Value
Pearson Correlation ($r$) 0.1210
Mean Absolute Error (MAE) 0.2512
Root Mean Square Error (RMSE) 0.3024
Cosine Similarity 0.6543
The results, presented in Table 3, show significant changes in the correlation metrics. The Pearson Correlation ($r$) dropped to 0.1210, indicating a poor linear relationship between the generated and reference mechanical properties. This suggests that without level 1 fine-tuning, the model struggles to capture the variations in the properties correctly. Additionally, the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) have increased, indicating larger discrepancies between the generated properties and the reference values. This highlights that the model is less accurate in predicting mechanical properties when fine-tuning is omitted. Furthermore, the Cosine Similarity decreased to 0.6543, suggesting a diminished directional similarity between the two sets of properties. This reduction reflects the model’s inability to maintain the overall trends and patterns of the reference properties without the fine-tuning step. These observations underscore the importance of level 1 fine-tuning in ensuring that the model performs optimally in predicting mechanical properties.

### v2

@@ 4.4 MaSp Motif-Property Correlation Analysis
The relationship between MaSp motifs and mechanical properties in spider dragline silk offers critical insights into how specific protein sequence patterns influence silk performance. Recent high-throughput studies have quantified key mechanical metrics -including toughness, Young’s modulus, tensile strength, and strain at break - of dragline silk from various spider species, identifying correlations between motifs in the repetitive core region of MaSps and variations in these properties (Arakawa et al., 2022).
@@
In previous studies, correlation analyses have primarily focused on the distinct motifs present in different MaSp types (MaSp1, MaSp2, and MaSp3) and their specific influences on mechanical properties (Craig et al., 2020; Nakamura et al., 2024a; Kono et al., 2021a; Arakawa et al., 2022). By contrast, the current work considers MaSp more abstractly, providing an overarching view of how various motifs may affect mechanical performance.
@@
Figure 4 presents the correlation heatmap between sequential features and mechanical properties. The sequential features include the count and coverage ($\frac{\text{motif size}\times\text{count}}{\text{sequence length}}$) of the specified motifs. The heatmap indicates that correlation values remain relatively low. While no strong correlations were observed, this visualization is still informative, demonstrating that individual motifs alone cannot fully explain the relationship between sequence variation and mechanical properties. This highlights the intricate and multi-factorial nature of silk mechanics, where the combined influence of multiple motifs and structural contexts collectively shapes the observed mechanical behavior.
@@
In the following, we present a row-wise analysis of the heatmap:
@@
Toughness is the total energy absorbed before failure. Higher toughness in dragline silk is associated with motifs that enhance tensile strength and/or extensibility. The correlation heatmap reveals that YGQGG count (+0.228) and coverage (+0.207) both show moderate positive correlations with toughness. Poly-Ala coverage also correlates positively with toughness, though slightly less strongly (+0.218). The AGQG coverage (-0.192) and SV count (-0.124) showed negative correlation with toughness .
@@
Young’s modulus represents the initial stiffness of silk fibers and is influenced by motifs that foster $\beta$-sheet crystallinity. For example, poly-Ala count correlates positively with Young’s modulus (+0.174). In our analysis, we also observed a positive impact of the count of YGQGG motif on Young’s modulus (+0.091), whereas the count of AGQG (-0.036) and SV (-0.139) motifs showed negative correlations.
@@
Tensile strength is the maximum stress that the silk fiber can withstand before failure. The correlation heatmap indicates that many of the same motifs influencing toughness also affect strength. The YGQGG motif count is positively correlated with tensile strength (+0.230) (Arakawa et al., 2022), making it one of the strongest sequence predictors of a stronger fiber in our analysis. Poly-Ala coverage (+0.140) and GPGXX coverage also positively correlate with tensile strength.
@@
Strain at break is the maximal elongation (expressed as a fraction of its original length) that the silk fiber can achieve before breaking. The correlations highlight AGQG coverage (–0.133) is the most striking correlation overall (and the only moderately strong negative value). Other features in this row hover near zero, indicating minimal correlation. The results also show impact of QQ motif on strain at break. Other features in this row hover near zero, indicating minimal or no correlation.
In summary, the correlation between mechanical properties and individual motifs in the generated sequences follows patterns reported in previous studies (Arakawa et al., 2022; Nakamura et al., 2024b). However, the overall weak correlations suggest that individual motifs alone cannot account for the observed variations in fiber mechanical properties (Arakawa et al., 2022). This study builds upon existing sequence-property relationships and leverages machine learning to better capture the complex interplay between sequential motifs and mechanical properties, enabling the design of novel MaSp sequences.
@@ 5.1.1 Conditional Generation Assessment: Evaluation on test dataset
To assess the self-consistency of the trained model, we evaluated its performance on the test sets from all five models trained during Stage 3 (Level 2 fine-tuning). Each model was tested on a set of 37 instances, resulting in a total of 185 unique test records. Each instance includes a MaSp repeat sequence along with its corresponding mechanical property set. The detailed composition of the test set is provided in the supplementary document SD2. The mechanical properties from these test instances were used as input conditions for sequence generation.
We note that the sequence–property relationship is not strictly one-to-one. The comparison is based on the hypothesis that similar mechanical properties can arise from sequences that share structural or motif-level features, particularly within the MaSp repeat regions. The objective was to qualitatively assess whether the generated sequences resemble natural ones associated with similar mechanical profiles.
To evaluate how well the model captures the relationship between sequence and function, we compared the generated sequences to their natural counterparts at both the sequence and structural levels. As shown in Figure 5, this dual-level analysis highlights the model’s ability to reproduce key biochemical and structural characteristics of spider silk proteins when conditioned on specific mechanical properties.
@@ 5.1.2 Novelty Assessment: BLAST Evaluation
To evaluate the novelty of the generated MaSp repeats, we performed BLAST analyses against a broader Spider Silkome spidroin repeat database. The search space consists of 11K naturally occurring spidroin sequences. It includes seven primary types of spider silk proteins major ampullate spidroins (MaSp), flagelliform spidroins (Flag), minor ampullate spidroins (MiSp), aggregate spidroins (AgSp), pyriform spidroins (PySp), aciniform spidroins (AcSp), and cylindriform/tubuliform spidroins (CySp) (Yarger et al., 2018a). A subset of seven generated sequences was compared with the two most closely related natural sequences, designated BLAST1 (top match) and BLAST2 (second top match) based on BLAST results. The selection was based on high query coverage, percent identity, and sequence length similarity to the generated sequences. This dataset is referred to as the "BLAST set", with details of all sequences and their sources provided in the supporting document SD3. All selected BLAST matches had an expect value (E) of $E\leq e^{-10}$, which ensured statistically significant similarities. Furthermore, we selected matches spanning different MaSp subtypes (MaSp1, MaSp2, MaSp3) to provide a comprehensive comparative analysis. The expect value (E) quantifies the probability of obtaining a match by chance in a given database size, decreasing exponentially as the alignment score (S) increases.
To assess uniqueness, we follow established criteria where sequences with similarity values below 50–60% are considered novel (Quan et al., 2023). The generated sequences meet this threshold when compared to broader protein databases, demonstrating the model’s ability to design distinct sequences that either do not exist in nature or have not yet been observed. Additionally, BLAST search results consistently classify the most similar existing sequences as MaSp, aligning with the intended spidroin type. This indicates that the generative modeling approach effectively captures the defining characteristics of specific spider silk proteins.
@@
Table 1: Quantitative comparison of mechanical property prediction accuracy between SpiderGPT (ours) and SilkomeGPT (baseline) based on trend similarity metrics computed over reference and generated mechanical property profiles. Bold values indicate better performance for each metric.
Metric SpiderGPT (ours) SilkomeGPT (baseline)
Pearson Correlation ($r$) 0.8884 0.8349
Spearman Correlation ($\rho$) 0.8343 0.7798
Mean Absolute Error (MAE) 0.0861 0.0963
Root Mean Square Error (RMSE) 0.1047 0.1162
$R^{2}$ Score 0.6383 0.5861
Cosine Similarity 0.9827 0.9783
@@ 5.2 Evaluation of Mechanical Property Prediction
Figure 7: Comparison of predicted and original mechanical properties for MaSp repeats in the test set. (a) Scatter plots show the correlation between predicted and actual values for normalized mechanical properties, with Pearson’s $r$ and ideal 1:1 (red dashed) line. (b) Kernel density plots compare natural and generated distributions for each property, including KS test $p$-values, means ($\mu$), and standard deviations ($\sigma$). Results indicate effective property prediction by the model.
To evaluate the model’s ability to estimate mechanical properties from MaSp repeat sequences, we use the test set of 185 instances described earlier. For each sequence in this set, we perform inference using the model’s reverse task—predicting mechanical properties from sequence input. The predicted properties are then compared against the corresponding ground-truth values.
We benchmark the performance of our model (SpiderGPT) against SilkomeGPT (Lu et al., 2024). Since SilkomeGPT does not focus exclusively on repeat regions, we use the original (unprocessed) sequences from the test set—that is, before repeat extraction—for both models to ensure a fair comparison in the mechanical property prediction task.
Prediction quality is assessed using five standard statistical metrics: Pearson’s correlation coefficient ($r$), Spearman’s rank correlation ($\rho$), mean absolute error (MAE), root mean square error (RMSE), and cosine similarity. The results, presented in Table 1, provide a quantitative comparison of the models’ ability to replicate the mechanical property patterns in the reference data. It is important to note that SilkomeGPT was trained on the entire Spider Silkome dataset, meaning that the test instances used here were part of its training data. Despite this advantage, SpiderGPT—trained only on repeat regions and without exposure to the test set—outperforms SilkomeGPT across all evaluated metrics.
The Pearson correlation coefficient was found to be $r=0.8884$, indicating a strong positive linear relationship between the predicted and reference values. This suggests that the model captures a substantial portion of the underlying variance in the data, even though the relationship is not perfectly linear. Additionally, the Spearman correlation coefficient was $\rho=0.8343$, reflecting a strong monotonic relationship. Unlike Pearson, Spearman correlation assesses the consistency of the rank order between two variables. A high $\rho$ value implies that when a reference property increases, the corresponding predicted property also tends to increase, even if the absolute values do not match precisely. This highlights the model’s ability to preserve the relative ordering of mechanical properties.
The MAE was measured to be 0.0861, indicating that, on average, predicted values differ only slightly from the reference values. The RMSE was slightly higher at 0.1047, consistent with its sensitivity to larger deviations. This implies that while most predictions are accurate, a few instances exhibit moderate error, which remains within acceptable bounds.
The coefficient of determination ($R^{2}$) was observed to be 0.6383. In simple linear regression, $R^{2}$ is equal to the square of the Pearson correlation coefficient, i.e., $R^{2}=r^{2}$ Substituting $r=0.89$, we obtain $R^{2}=(0.89)^{2}\approx 0.79$. However, the current dataset involves multiple output dimensions and includes non-linear dispersion, which can lower the observed $R^{2}$ despite a high average correlation. This explains the discrepancy between the theoretical and observed values, with $R^{2}=0.6383$ being a reasonable outcome in this context.
In addition, the cosine similarity between predicted and original property vectors was calculated to be 0.9827. This high value indicates strong directional alignment, implying that the overall structural trends are well preserved even if individual values are not perfectly matched.
A property-level analysis is provided in Figure 7. Panel (a) displays scatter plots comparing predicted and original values across the four mechanical properties: toughness, Young’s modulus, tensile strength, and strain at break. Each subplot includes the Pearson correlation coefficient ($r$) and the coefficient of determination ($R^{2}$), offering a detailed view of model performance. All properties exhibit strong positive correlations, with tensile strength achieving the highest $R^{2}$ (0.793) and toughness the lowest (0.689), indicating robust but slightly variable predictive accuracy across property types.
Panel (b) of Figure 7 shows kernel density estimates comparing the distributions of natural and generated property values. Each subplot includes the Kolmogorov–Smirnov (K–S) test $p$-value, along with the mean ($\mu$) and standard deviation ($\sigma$) for both distributions. High $p$-values across all properties (e.g., 0.752 for strain at break) suggest that the generated properties are statistically consistent with the natural ones, as measured by the K–S test (Massey Jr, 1951).
@@ 5.3.2 Without first level fine tuning
ProtGPT2 is a decoder-only transformer model that has been pre-trained on the protein space using the UniRef50 database (version 2021_04) (The UniProt Consortium, 2024), which contains 10 million protein instances. However, the Sequence-to-Mechanical Property Correlation dataset, curated from the Spider Silkome database, contains only 592 instances. This small dataset is insufficient for training the model to generate MaSp repeat sequences while simultaneously learning the correlation between sequences and mechanical properties. To address this limitation, we split the training into two stages.
@@
To further investigate these discrepancies, we compare the structural predictions of sequences generated by a model trained with only one fine-tuning stage against those produced by a model trained with both levels of fine-tuning. Figure 9 illustrates the molecular configurations of the generated sequences, revealing structural differences arising from the absence of MaSp-specific fine-tuning. Structural predictions were obtained using Omegafold (Wu et al., 2022), providing insights into the conformational tendencies of the sequences. Notably, the generated sequences without level 1 fine-tuning exhibited helical structures, which are typically associated with non repeat regions like terminal domains (NTD/CTD). These unwanted elements further highlight the necessity of level 1 fine-tuning in ensuring the correct generation of MaSp repeats.
@@
Table 3: Comparison of generated and reference properties without level 1 finetuning
Metric Mean Value
Pearson Correlation ($r$) 0.3628
Spearman Correlation ($\rho$) 0.3301
Mean Absolute Error (MAE) 0.2911
Root Mean Square Error (RMSE) 0.3447
R square ($r^{2}$) -2.7196
Cosine Similarity 0.8255
Here, we analyze the impact of skipping Level 1 fine-tuning on the reverse task of mechanical property estimation. Using the same test set, we prompted the model to predict mechanical properties from sequence inputs and compared the predicted values against their ground truth counterparts. To assess performance, we compared trend similarity metrics with and without Level 1 fine-tuning. When the model was trained without the 6,000 MaSp repeat sequences used in Level 1, its ability to recover accurate property trends declined significantly, as shown in Table 3. The mean Pearson correlation dropped from $0.8884$ to $0.3628$, and the mean Spearman correlation fell from $0.8343$ to $0.3301$, indicating a considerable loss in both linear and rank-order consistency.
Prediction errors also increased substantially: MAE rose from $0.0861$ to $0.2911$, and RMSE increased from $0.1047$ to $0.3447$. Most notably, the $R^{2}$ score deteriorated from $0.6383$ to 2.7196-2.7196, suggesting that without Level 1 fine-tuning, the model performs worse than a simple mean-based predictor. Cosine similarity also decreased from $0.9827$ to $0.8255$, reflecting weakened directional agreement between predicted and reference property vectors.
These results highlight the essential role of Level 1 fine-tuning in learning robust sequence embeddings that generalize well to downstream property prediction. Its absence leads to diminished trend fidelity and overall performance degradation.